OpenStudy (anonymous):

Solve for the variable in this equation. 1/2y-2/3=7/8

OpenStudy (anonymous):

I believe you mean "1/(2y) - 2/3 = 7/8" First solve for 1/(2y). What do you get?

OpenStudy (anonymous):

(1/2y)-(2/3)=7/8 denominators are 2y, 3, 8 LCD of them is 24y so multiply 24y to both sides 12-16y=21y 12-16y+16y=21y+16y 12=37y 37y/37=12/37 y=12/37

OpenStudy (anonymous):

Thanks:)